What is the mass of a car moving forward with a force of 3000N and an acceleration of 0.15 m/s^2?
1
Expert's answer
2018-09-10T12:58:08-0400
We can find the mass of the car from the Newton’s Second Law of Motion:
F=ma,
m=F/a=(3000 N)/(0.15 m/s^2 )=20000 kg.